A particle is located at the vector position  = (2 + 4) m and the force acting on it is  = 4 + 3 N. What is the torque about the origin ?
Answer #1

The torque on a particle due to the force \vec{F} , and at a position \vec{r} is given by

  \vec{\tau } = \vec{r} \times \vec{F}

Given,

  \vec{r} = 2 \,\hat{i} + 4 \,\hat{j} (in m) and \vec{F} = 4 \,\hat{i} + 3 \,\hat{j} (in N)

Therefore, torque on the particle about the origin is

  \vec{\tau} =( 2 \,\hat{i} + 4 \,\hat{j}) \times ( 4 \,\hat{i} + 3 \,\hat{j})

or,   \vec{\tau} = \begin{vmatrix} \vec{i} & \vec{j} & \vec{k} \\ 2 &4 &0 \\ 4& 3 &0 \end{vmatrix}

or, \vec{\tau} =\vec{i} . (0-0) - \vec{j } . (0-0)+\vec{k} . (6-16)

or,   \vec{\tau} =- 10 \,\vec{k} (in N.m)